# Balancing Specialization and Adaptation in a Transforming Scientific Landscape

### 03 Jan 2024 v2

@@ 2.4 Modelling trajectories
Physicists’ trajectories are modeled with a hierarchical multinomial logistic model. The intuition underlying the model is illustrated in Figures 1(a),1(b),1(c), and a more formal representation of the model is given in Figure 1(d). The occurrences of keywords that belong to each research are $k$ in papers by $a$ in the late time period, $\bm{Y_{a}}\in\mathbb{R}^{K}$, are assumed to be a function of the initial research portfolio ($\bm{X_{a}}$), thus capturing the intuition illustrated in Figure 1 that research agendas are transformed via strategic transfers of knowledge from one research area to another. $\bm{Y_{a}}$ is also a function of a mixing matrix $\theta_{a}$ that measures the extent to which attention has been redistributed to different research areas. $\theta_{a}$ is drawn from a hierarchical process, thus capturing the “average” behavior of the cohort. It is also a function of covariates ($\bm{Z_{a}}$ in Figure 1(d)), in this case intellectual and social capital, which are hypothesized to affect how individuals have revised their research agenda.
Formally speaking, we assume that $\bm{Y_{a}}$ derives from a multinomial process involving linear combinations of $(x_{ak})$:
@@
Our model is strongly connected to Optimal Transport [49, 50]. Optimal Transport seeks the most efficient way to “transport” a source distribution (say, $\bm{x}=(x_{k})$) to a target distribution (e.g., $\bm{y}=(y_{k^{\prime}}$)) through “transfers” $(\theta_{kk^{\prime}})$ across their components, given a cost function $\sum_{k,k^{\prime}}x_{k}\theta_{kk^{\prime}}C_{kk^{\prime}}$ (where $C_{kk^{\prime}}$ is a cost matrix) [51]. The difference is that our Bayesian approach estimates transfers $\theta_{kk^{\prime}}$ by minimizing a likelihood rather than a cost function. The connection with Optimal Transport suggests an “economic” interpretation of the reallocation of research efforts, which will be leveraged in Section 3.2 to show that patterns of change in research interests are mainly structured by learning costs.
@@ 3.1 Transfers across research areas
Figure 3 shows the aggregate transfers of attention at the level of the cohort unveiled by the model. The width of each flow is proportional to $\sum_{a}X_{ak}\theta_{akk^{\prime}}$ (summing over each author $a$). The most obvious feature is the overall conservatism of physicists toward their research area, i.e. specialization. Certain research areas have significantly grown, most notably “dark matter” and “AdS/CFT”, while others – “neutrinos and flavor physics” and “string theory and supergravity” – have significantly shrunk. The increasing interest in dark matter seems to be mainly fueled by a shift away from “neutrinos and flavor physics” and “electroweak sector”, while the growth in AdS/CFT and black hole research has been at the expense of “string theory and supergravity” research [^12] This converges with physicist Peter Woit’s controversial assessment that “string theorists” are no longer doing string theory per se, though they keep identifying themselves as string theorists. As Peter Woit puts it, citing the 2022 “Strings” conference: “one thing that stands out is that the string theory community has almost completely stopped doing string theory.”; and, “[presentations’ titles] make very clear what the string theory community has found to replace string theory: black holes” (Woit, 2022, https://www.math.columbia.edu/~woit/wordpress/?p=12981);. Smaller research areas such as “classical fields”, “dynamics” or “thermodynamics” seem less cohesive, i.e., they are subject to larger transfers. These bodies of knowledge are linguistically coherent (and therefore appear as topics of their own) but yet are not able to sustain careers over the long run by themselves.
@@ 3.2 The structuring effect of learning costs on scientists’ behavior
As research priorities change, research efforts must be efficiently reallocated among scientists, given their prior knowledge. Here, we leverage the similarity between our model of scientists’ trajectories and Optimal Transport [49, 50] to formulate the observed behavior in economic terms and demonstrate that the shifts of attention are structured by the minimization of “learning costs”.
@@
Let $\bm{x}=\sum_{a}\bm{x_{a}}$ be the cohort’s initial distribution of attention across research areas (summing over each author $a$), and $\bm{y}=\sum_{a}\bm{y_{a}}$ the late distribution. Let us further assume that $\bm{x}$ and $\bm{y}$ can be considered “fixed” by the institutions (scientific leadership, laboratories, funding agencies, etc.) that define scientific priorities throughout time. In order to achieve the distribution of research efforts $\bm{y}$ given the previous distribution $\bm{x}$, some scientists must redirect their attention away from certain research areas (for which $y_{k}<x_{k}$) and towards more pressing ones (for which $y_{k}>x_{k}$). What is the most efficient way to reallocate research efforts and achieve the transition from $\bm{x}$ to $\bm{y}$? Intuitively, research areas should be assigned to scientists in a way that requires as few of them as possible to acquire new knowledge – in other words, in a way that minimizes learning costs, given the way knowledge is distributed among individuals. This, as we show, can be framed as an Optimal Transport problem. Let $T_{kk^{\prime}}$ be the coupling matrix that encodes how much attention has been shifted from one research area $k$ to a research area $k^{\prime}$. There are two constraints on $T_{kk^{\prime}}$: $\sum_{k^{\prime}}T_{kk^{\prime}}=x_{k}$, and $\sum_{k}T_{kk^{\prime}}=y_{k^{\prime}}$. These constraints encode the need to adapt (since $\bm{y}\neq\bm{x}$). But $T$ must also minimize learning costs ($C$), which are linear in $T$: $C=\sum_{k,k^{\prime}}T_{kk^{\prime}}C_{kk^{\prime}}$, where $C_{kk^{\prime}}$ is a cost matrix. The problem of finding the couplings $T_{kk^{\prime}}$ that minimize the “transportation” costs (given a cost matrix $C_{kk^{\prime}}$ and the constraints on $T_{kk^{\prime}}$) is an instance of Optimal Transport problem [51]. Examples of Optimal Transport problems include how to efficiently transport ore from a set of mines to a set of factories given the distances between them, or the optimal assignment of workers to firms [52]. In our case, the couplings are known (they have been recovered by our model from Section 2.4), and we would like to infer the underlying cost matrix that these couplings minimize. The transfers from a research area $k$ to $k^{\prime}$ for each individual are simply $x_{ak}\theta_{akk^{\prime}}$. Summing over individuals yields the coupling matrix $T_{kk^{\prime}}=\sum_{a}x_{ak}\theta_{akk^{\prime}}$, which measures how much attention was shifted away from $k$ and toward $k^{\prime}$ at the cohort level.
[…]
@@
Under changing circumstances, research efforts must be reallocated efficiently. Therefore, scientific norms and institutions address an Optimal Transport problem by providing incentives for scientists to conform to new research imperatives, in a way that minimizes learning costs. While this observation characterizes the cohort behavior, drivers of differences among individuals are considered next.
@@ 3.3.1 Magnitude of change and capital
$c_{a}$ is comprised between 0 (if research attention has remained identically distributed) and 1 (if the research agenda has been entirely redistributed). The distribution of $c_{a}$ (Figure 5) shows that large values are rare, with 50% of authors lying between 0.21 and 0.40. However, the absolute value of $c_{a}$ has limited interpretability, given that it depends on the choice of “cognitive scope” (the broadness of research areas) and the duration of the time periods considered [13]. Moreover, $c_{a}$ is never expected to be exactly zero due to random fluctuations.
@@
Unsurprisingly, both research areas that have shrunk considerably have a significant positive effect on migration score. “Collider physics”, on the other hand, has a negative effect on the magnitude of change. This is the case ceteris paribus; in other words, physicists whose primary category is “Collider physics” are more conservative than other physicists with similar capital diversity, power, and position stability. In fact, these physicists also have less diverse capital, and are by far the most conservative, with an average change score 24% lower than the rest of the cohort. Indeed, the long time-scales of collider experiments provide stable opportunities to physicists in that area [55, p. 138], thus promoting conservatism. The variance explained remains low ($R^{2}=0.15$), and all these factors only partially explain differences between individuals.
@@ 3.3.2 Diversification of research interests versus consolidation of expertise
The diversity of intellectual capital has a strong positive effect on the probability of exiting a research area; intuitively, scientists with diverse interests can afford to disengage from certain research areas. Excess diversity of social capital increases the probability of entering new research areas, but has no discernible effect on the probability of exiting research areas. In contrast, “power” significantly decreases the probability of leaving a research area. This suggests that having many collaborators allows scientists to remain committed to many research areas with minimal personal investment, thus stabilizing their research agendas (as previously observed). Moreover, “power” also decreases the probability to of entering new research areas; this suggests that powerful scientists have less incentives to invest resources in new topics. String theory & supergravity has a significant positive effect of both entering and exiting research areas (suggesting it has been “replaced” by another topic by many physicists). “Collider physics”, however, significantly decreases the probability of entering new areas. Indeed, the LHC has secured this research area for the decades to come: for instance, the HL-LHC project (High-Luminosity LHC), which consists in upgrading the LHC in order to higher collision rates, is scheduled to start in 2029 at the time of writing [^17] See https://home.cern/science/accelerators/high-luminosity-lhc. Note that until a successor to the LHC is under active development, its prolongation can help preserve the know-how necessary for the design and operation of such complex experiments.. “Dark matter” has a negative effect on the probability of entering a new research area, possibly because scientists with prior commitment to this research area had less incentive to diversify their research portfolio and more incentive to focus increasingly on this topic given its success over the following years. Again, these conclusions hold as an alternative measure of diversity is considered (A.7.3, Table 4) and if “brokerage” is used as a measure of power.
@@ 3.3.3 Why diversity promotes change
Access to diverse cognitive resources promotes change. To see why, it is insightful to look into how the concentration of scientists’ intellectual capital in each research area $k$ ($\bm{I_{a}}=(I_{ak})$) affects their trajectories. In the model introduced in Section 2.4, the diagonal coefficients of the $\gamma$ matrix measure the effect of having intellectual resources in a certain area on the commitment to this area. The matrix is shown in Figure 7(a). Most coefficients on the diagonal are significantly positive, thus suggesting that physicists with a strong specialization in a research area have tend to further exploit and consolidate their specialization in this area.
The effect of social capital on transfers from one research area to another, given by the matrix $\delta_{kk^{\prime}}$ (see equation 6), is shown in Figure 7(b). Statistically significant effects are always positive: having collaborators with expertise in a target area generally increases the likelihood of redirecting attention toward this area.
@@ 4 Discussion
Then, leveraging the connection between our model and Optimal Transport (OT), we showed that the reallocation of research efforts among scientists is structured by learning costs. Indeed, under changing circumstances, scientific institutions must address an OT problem by efficiently reallocating research efforts in a way that balances learning costs and the imperative to adapt to new circumstances. This has the effect of enhancing the utility of the scientific capital disseminated amongst scientists as the perceived payoff of certain research areas increases. The connection with OT is methodologically useful: Inverse Optimal Transport allows one to derive cost functions from observed behavior, and thus offers a potential way to better connect empirical data with formal rational-choice models of scientists’ behavior. Moreover, it has the potential of informing policy by identifying potential bottlenecks if research efforts were to be reallocated in certain ways (as one can use OT to estimate the “cost” of various counterfactual scenarios).
@@
The longitudinal comparative analysis of physicists’ trajectories revealed that the diversity of intellectual and social capital is positively associated with change: diversity promotes adaptability under new circumstances, and therefore diversifying research portfolios is a reasonable strategy when the future is uncertain. However, enforcing diversity can lead to suboptimal allocations of research efforts under stable circumstances [25]. Differences among research areas are found: physicists expert in particle colliders have remained particularly conservative, possibly because they have secured long-term research opportunities (thanks to very large investments in particle accelerators like the LHC). There is also evidence that physicists specialized in dark matter have been consolidating their specialization, presumably because the increasing popularity of the topic encouraged them to double down their investment in this research area. Higher concentrations of intellectual capital in certain research areas generate stronger commitment to these research areas; therefore, specialized scientists are more at risk of being trapped in a sunk cost fallacy as their expertise becomes unsuitable for new circumstances. However, specialized scientists can offset the risks associated with specialization by diversifying their social network. This raises the possibility of free-riding, as scientists are encouraged to focus on what seems most promising at the time and let their peers take the risk of exploring alternatives until their value is established [56]. Additionally, social capital plays an increasingly important role as scientists expand their research agenda further beyond their specialization, as observed by [7], suggesting that collaborations are crucial in overcoming cognitive barriers between research areas. Unlike diversity, “power” (that is, the magnitude of social capital) and institutional stability are associated with more stability and conservatism in research interests. We suggest that coalitions and institutional resources can safeguard individuals from pressures to adapt. The negative association between change and institutional stability, albeit weak ($\mu(\beta)=-0.14$), suggests that scientific institutions may exploit temporary contracts to compell the scientific workforce to adapt.
@@
Throughout the paper, various methodological limitations were raised. Most of them pertain to data selection. For example, the requirement on the quantity of authors’ publications makes the cohort “unrepresentative” and implies a survivorship bias. A second issue, already noted by [13, ], lies in the arbitrariness in the choice of cognitive scope. Previous works (e.g. [4, 5, 14, 7]) have typically relied on the hierarchical PACS classification of physics literature, such that cognitive scales were imposed. For this work, the literature was divided into 16 topics that captured the features of change in the epistemic landscape discussed above, i.e. the rise of new probes of the cosmos, but other scales could have been considered. However, the topic model recovers a flat structure that does not explain the correlations between different research areas. A hierarchical structure of knowledge might be more suitable for capturing what allows physicists to switch from one research area to another.

### 14 Jan 2024 v3

@@ 2.4 Modelling trajectories
Physicists’ trajectories are modeled with a hierarchical multinomial logistic model. The intuition underlying the model is illustrated in Figures 1(a),1(b),1(c), and a more formal representation of the model is given in Figure 1(d). The occurrences of keywords that belong to each research are $k$ in papers by $a$ in the late time period, $\bm{Y_{a}}\in\mathbb{R}^{K}$, are assumed to be a function of the initial research portfolio ($\bm{X_{a}}$), thus capturing the intuition illustrated in Figure 1 that research agendas are transformed via strategic transfers of knowledge from one research area to another. $\bm{Y_{a}}$ is also a function of a mixing matrix $\theta_{a}$ that measures the extent to which attention has been redistributed to different research areas. $\theta_{a}$ is drawn from a hierarchical process, thus capturing the “average” behavior of the cohort. It is also a function of covariates ($\bm{Z_{a}}$ in Figure 1(d)), in this case intellectual and social capital, which are hypothesized to affect individual behaviors. Formally speaking, $\bm{Y_{a}}$ is assumed to derive from a multinomial process involving linear combinations of $(x_{ak})$:
@@
Our model is strongly connected to Optimal Transport [48, 49]. Optimal Transport seeks the most efficient way to “transport” a source distribution (say, $\bm{x}=(x_{k})$) to a target distribution (e.g., $\bm{y}=(y_{k^{\prime}}$)) through “transfers” $(\theta_{kk^{\prime}})$ across their components, given a cost function $\sum_{k,k^{\prime}}x_{k}\theta_{kk^{\prime}}C_{kk^{\prime}}$ (where $C_{kk^{\prime}}$ is a cost matrix) [50]. The difference is that our Bayesian approach estimates transfers $\theta_{kk^{\prime}}$ by minimizing a likelihood rather than a cost function. The connection with Optimal Transport suggests an economic interpretation of the reallocation of research efforts, which will be leveraged in Section 3.2 to show that patterns of change in research interests are shaped by learning costs.
@@ 3.1 Transfers across research areas
Figure 3 shows the aggregate transfers of attention at the level of the cohort unveiled by the model. The width of each flow is proportional to $\sum_{a}X_{ak}\theta_{akk^{\prime}}$ (summing over each author $a$). The most obvious feature is the overall conservatism of physicists toward their research area, i.e. specialization. HEP has accumulated large amounts of human and economic capital – thus securing certain research areas for a long time – and is associated with more conservative research interests compared to other areas of physics [5]. Certain research areas have significantly grown, most notably “dark matter” and “AdS/CFT”, while others – “neutrinos and flavor physics” and “string theory and supergravity” – have significantly shrunk. The increasing interest in dark matter seems to be mainly fueled by a shift away from “neutrinos and flavor physics” and “electroweak sector”, while the growth in AdS/CFT and black hole research has been at the expense of “string theory and supergravity” research [^12] This converges with physicist Peter Woit’s controversial assessment that “string theorists” are no longer doing string theory per se, though they keep identifying themselves as string theorists. As Peter Woit puts it, citing the 2022 “Strings” conference: “one thing that stands out is that the string theory community has almost completely stopped doing string theory.”; and, “[presentations’ titles] make very clear what the string theory community has found to replace string theory: black holes” (Woit, 2022, https://www.math.columbia.edu/~woit/wordpress/?p=12981);. Smaller research areas such as “classical fields”, “dynamics” or “thermodynamics” seem less cohesive, i.e., they are subject to larger transfers. These bodies of knowledge are linguistically coherent (and therefore appear as topics of their own) but yet are not able to sustain careers over the long run by themselves.
@@ 3.2 The structuring effect of learning costs on scientists’ behavior
As research priorities change, research efforts must be efficiently reallocated among scientists, given their prior knowledge. Here, we leverage the similarity between our model of scientists’ trajectories and Optimal Transport (OT) [48, 49] to formulate the observed behavior in economic terms and demonstrate that the shifts of attention are structured by the minimization of “learning costs”.
@@
Let $\bm{x}=\sum_{a}\bm{x_{a}}$ be the cohort’s initial distribution of attention across research areas (summing over each author $a$), and $\bm{y}=\sum_{a}\bm{y_{a}}$ the late distribution. Let us further assume that $\bm{x}$ and $\bm{y}$ can be considered “fixed” by the institutions (scientific leadership, laboratories, funding agencies, etc.) that define scientific priorities throughout time. In order to achieve the distribution of research efforts $\bm{y}$ given the previous distribution $\bm{x}$, some scientists must redirect their attention away from certain research areas (for which $y_{k}<x_{k}$) and towards more pressing ones (for which $y_{k}>x_{k}$). What is the most efficient way to reallocate research efforts and achieve the transition from $\bm{x}$ to $\bm{y}$? Intuitively, research areas should be assigned to scientists in a way that requires as few of them as possible to acquire new knowledge – in other words, in a way that minimizes learning costs, given the way knowledge is distributed among individuals. This, as we show, can be framed as an Optimal Transport problem. Let $T_{kk^{\prime}}$ be the coupling matrix that encodes how much attention has been shifted from one research area $k$ to a research area $k^{\prime}$. There are two constraints on $T_{kk^{\prime}}$: $\sum_{k^{\prime}}T_{kk^{\prime}}=x_{k}$, and $\sum_{k}T_{kk^{\prime}}=y_{k^{\prime}}$. These constraints encode the need to adapt (since $\bm{y}\neq\bm{x}$). But $T$ must also minimize learning costs ($C$), which are linear in $T$: $C=\sum_{k,k^{\prime}}T_{kk^{\prime}}C_{kk^{\prime}}$, where $C_{kk^{\prime}}$ is a cost matrix. The problem of finding the couplings $T_{kk^{\prime}}$ that minimize the “transportation” costs (given a cost matrix $C_{kk^{\prime}}$ and the constraints on $T_{kk^{\prime}}$) is an instance of Optimal Transport problem [50]. Examples of Optimal Transport problems include how to efficiently transport ore from a set of mines to a set of factories given the distances between them, or the optimal assignment of workers to firms [51]. In our case, the couplings are known (they have been recovered by the model from Section 2.4), and we would like to infer the underlying cost matrix that these couplings minimize. The transfers from a research area $k$ to $k^{\prime}$ for each individual are simply $x_{ak}\theta_{akk^{\prime}}$. Summing over individuals yields the coupling matrix $T_{kk^{\prime}}=\sum_{a}x_{ak}\theta_{akk^{\prime}}$, which measures how much attention was shifted away from $k$ and toward $k^{\prime}$ at the cohort level.
The problem of recovering the cost matrix $C_{kk^{\prime}}$ given the couplings is called the Inverse Optimal Transport problem. Below, this problem is solved using a method proposed and tested on European migration patterns in [52]. This method requires to put a prior on $P(C_{kk^{\prime}})$ – indeed, there are infinitely many cost matrices that can yield the same optimum, and priors are needed to lift this degeneracy. Following [52], we consider a prior such that $\sum_{k,k^{\prime}}C_{kk^{\prime}}=\mathrm{cst}$. We assume that $\mathbb{E}(C_{kk^{\prime}})\propto\mathrm{softmax}(\beta\times(1-\nu_{kk^{\prime}}))$, where $\nu_{kk^{\prime}}$ is the fraction of physicists who already held expertise in $k^{\prime}$ among those who already held expertise in $k$, and $\beta\sim\mathcal{N}(0,1)$ is the effect of learning costs on $C_{kk}$. If $\nu_{kk^{\prime}}\sim 1$ (i.e. $1-\nu_{kk^{\prime}}\sim 0$), then shifting attention from $k$ to $k^{\prime}$ does not entail the acquisition of additional knowledge, and $C_{kk^{\prime}}$ should be almost zero. If $\nu_{kk^{\prime}}=0$ (i.e. $11-\nu_{kk^{\prime}}=$ 1), any scientist shifting from $k$ toward $k^{\prime}$ must acquire new knowledge, and the cost should be maximal. If actual behaviors do involve the minimization of learning costs, we should observe a strong correlation between $C_{kk^{\prime}}$ and the “knowledge gap”, which is $1-\nu_{kk^{\prime}}$.
Using the “MetroMC” algorithm proposed in [52], we empirically recover $C_{kk^{\prime}}$, the underlying cost matrix (Appendix A.6, Figure 14). We find a strong correlation with the “knowledge gap”: $R(C_{kk^{\prime}},1-\nu_{kk^{\prime}})=0.72$, such that $R^{2}=0.52$ (see Appendix A.6, Figure 14(a)). The empirical cost of shifting research efforts from one research area to another is therefore shaped by learning costs. The derivation of $C_{kk^{\prime}}$ is potentially useful: one could in principle predict aggregate transfers of attention given a counterfactual target distribution of research efforts $\bm{y}$ using optimal transport and plugging-in $C_{kk^{\prime}}$ as the cost matrix.
@@
Under changing circumstances, research efforts must be reallocated efficiently. Therefore, scientific norms and institutions address an Optimal Transport problem by providing incentives for scientists to conform to new research imperatives, in a way that factors “learning costs”. While this observation characterizes the cohort behavior, drivers of differences among individuals are considered next.
@@ 3.3.1 Magnitude of change and capital
$c_{a}$ is comprised between 0 (if research attention has remained identically distributed) and 1 (if the research agenda has been entirely redistributed). The distribution of $c_{a}$ (Figure 5) shows that large values are rare, with 50% of authors lying between 0.21 and 0.40. However, the absolute value of $c_{a}$ has limited interpretability, given that it depends on the choice of “cognitive scope” (the broadness of research areas) and the duration of the time periods considered [13]. Moreover, $c_{a}$ is never expected to be exactly zero due to random fluctuations and measurement noise.
@@
Unsurprisingly, both research areas that have shrunk considerably have a significant positive effect on migration scores. “Collider physics”, on the other hand, has a negative effect on the magnitude of change ceteris paribus. All effects combined, physicists whose primary category is “Collider physics” are by far the most conservative, with an average change score 24% lower than the rest of the cohort. Indeed, the long time-scales of collider experiments provide stable opportunities to physicists in that area [53, p. 138], thus promoting conservatism. The variance explained remains low ($R^{2}=0.15$), and all these factors only partially explain differences between individuals.
@@ 3.3.2 Diversification of research interests versus consolidation of expertise
The diversity of intellectual capital has a strong positive effect on the probability of exiting a research area; intuitively, scientists with diverse interests can afford to disengage from certain research areas. Excess diversity of social capital increases the probability of entering new research areas, but has no discernible effect on the probability of exiting research areas. In contrast, “power” significantly decreases the probability of leaving a research area. This suggests that having many collaborators allows scientists to remain committed to many research areas with minimal personal investment, thus stabilizing their research agendas (as previously observed). Moreover, “power” also decreases the probability to of entering new research areas; this suggests that powerful scientists have less incentives to invest resources in new topics. String theory & supergravity has a significant positive effect of both entering and exiting research areas (suggesting it has been “replaced” by another topic by many physicists). “Collider physics”, however, significantly decreases the probability of entering new areas. Indeed, the LHC has secured this research area for decades, and the High-Luminosity LHC, an upgraded of the LHC set to achieve higher collision rates, is only scheduled to start in 2029161616See https://home.cern/science/accelerators/high-luminosity-lhc. Note that until a successor to the LHC is under active development, its prolongation can help preserve the know-how necessary for the design and operation of such complex experiments.. “Dark matter” has a negative effect on the probability of entering a new research area, possibly because scientists with prior commitment to this research area had less incentive to diversify their research portfolio and more incentive to focus increasingly on this topic given its success over the following years. Again, these conclusions hold as an alternative measure of diversity is considered (A.7.3, Table 4) and if “brokerage” is used as a measure of power.
@@ 3.3.3 Why diversity promotes change
Access to diverse cognitive resources promotes change. To see why, it is insightful to look into how the concentration of scientists’ intellectual capital in each research area $k$ ($\bm{I_{a}}=(I_{ak})$) affects their trajectories. In the model introduced in Section 2.4, the diagonal coefficients of the $\gamma$ matrix measure the effect of having intellectual resources in a certain area on the commitment to this area. The matrix is shown in Figure 7(a). Most coefficients on the diagonal are significantly positive, thus suggesting that physicists with a strong specialization in a research area tend to further exploit and consolidate their specialization in this area.
The effect of social capital on transfers from one research area to another, given by the matrix $\delta_{kk^{\prime}}$ (see equation 6), is shown in Figure 7(b). Statistically significant effects are always positive: having moreand more collaborators with expertise in a target area generally increases the likelihood of redirecting attention toward this area, in line with a very recent observation that “the probability for a scholar to start working on a new topic increases with the number of previous collaborators” [54].
@@ 4 Discussion
Then, leveraging the connection between our model and Optimal Transport (OT), we showed that the reallocation of research efforts among scientists is shaped by learning costs. Indeed, under changing circumstances, scientific institutions must address an OT problem by efficiently reallocating research efforts in a way that balances learning costs and the imperative to adapt to new circumstances. This has the effect of enhancing the utility of the scientific capital disseminated amongst scientists as the perceived payoff of certain research areas change. The connection with OT is methodologically useful: Inverse Optimal Transport allows one to derive cost functions from observed behavior, and thus offers a potential way to better connect empirical data with formal rational-choice models of scientists’ behavior that postulate utility and cost functions. Moreover, it has the potential of informing policy by identifying potential bottlenecks if research efforts were to be reallocated in certain ways (as one can use OT to estimate the “cost” of various counterfactual scenarios).
@@
The longitudinal comparative analysis of physicists’ trajectories revealed that the diversity of intellectual and social capital is positively associated with change: diversity promotes adaptability under new circumstances, and therefore diversifying research portfolios is a reasonable strategy when the future is uncertain. However, enforcing diversity can lead to suboptimal allocations of research efforts under stable circumstances [25]. Differences among research areas are found: physicists expert in particle colliders have remained particularly conservative, possibly because they have secured long-term research opportunities (thanks to very large investments in particle accelerators like the LHC). There is also evidence that physicists specialized in dark matter have been consolidating their specialization, presumably because the increasing popularity of the topic encouraged them to double down their investment in this research area. Higher concentrations of intellectual capital in certain research areas generate stronger commitment to these research areas; therefore, specialized scientists are more at risk of being trapped in a sunk cost fallacy as their expertise becomes unsuitable for new circumstances. However, specialized scientists can offset the risks associated with specialization by diversifying their social network. This raises the possibility of free-riding, as scientists are encouraged to focus on what seems most promising at the time and let their peers take the risk of exploring alternatives until their value is established [55]. Additionally, social capital plays an increasingly important role as scientists expand their research agenda further beyond their specialization, as observed by [7], suggesting that collaborations are crucial in overcoming cognitive barriers between research areas. Unlike diversity, “power” (that is, the magnitude of social capital) and institutional stability are associated with more stability and conservatism in research interests. We suggest that coalitions and institutional resources can safeguard individuals from pressures to adapt. The negative association between change and institutional stability, albeit weak ($\mu(\beta)=-0.14$), suggests that scientific institutions may be exploiting temporary contracts to compel the scientific workforce to adapt.
@@
Throughout the paper, various methodological limitations were raised. Most of them pertain to data selection. For example, the requirement on the quantity of authors’ publications makes the cohort “unrepresentative” and implies a survivorship bias. A second issue, already noted by [13, ], lies in the arbitrariness in the choice of cognitive scope. Previous works (e.g. [4, 5, 7]) have typically relied on the hierarchical PACS classification of physics literature, such that cognitive scales were imposed. For this work, the literature was divided into 16 topics that captured the features of change in the epistemic landscape discussed above, i.e. the rise of new probes of the cosmos, but other scales could have been considered. However, the topic model recovers a flat structure that does not explain the correlations between different research areas, and a hierarchical structure of knowledge might be more suitable for capturing what allows physicists to switch from one research area to another. Moreover, although the unsupervised topic model approach is arguably a better proxy to linguistic and cognitive behavior, it introduces noise which could in part explain the low predictive power of the models of change in scientists’ research agendas.